# Nonreciprocal superradiant quantum phase transition induced by magnon Kerr effect

### v1

@@ III Nonreciprocal superradiant quantum phase transition
IV Conclusion
In summary, we have demonstrated that the MKE in a cavity magnonic system offers a viable pathway to achieve a nonreciprocal SQPT Chiacchio23 ; Zhu24 ; Xu24 . The proposed setup, comprising a YIG sphere coupled to a microwave cavity, exhibits markedly different phase diagrams under positive and negative Kerr coefficients-controlled by orienting the bias magnetic field along different crystallographic axes of the YG sphere Wang18 ; Zhang19 . These distinct phase boundaries directly lead to nonreciprocal critical behavior, as confirmed by analyzing the magnon occupation and its fluctuations around the transition. The introduction of a bidirectional contrast ratio further allows quantitative characterization of the nonreciprocity. Our results establish cavity magnonics as a promising platform for exploring nonreciprocal SQPT and may inform future designs of quantum devices requiring directional control.
@@ Acknowledgments
This work is supported by the National Natural Science Foundation of China (Grants No. 12205069 and No. 12204139), the National Key Research and Development Program of China (Grant No. 2024YFA1408900), the Natural Science Foundation of Zhejiang Province (Grant No. LY24A040004), Zhejiang Province Key R&D Program of China (Grant No. 2025C01028), and Shenzhen International Quantum Academy (Grant No. SIQA2024KFKT010).
@@
where $a^{{\dagger}}$ and $m^{{\dagger}}$ ($a$ and $m$) are the creation (annihilation) operators for the cavity and magnon modes with frequencies $\omega_{a}$ and $\omega_{m}$, respectively, $K$ is the Kerr coefficient, $g_{m}$ is the coupling strength between cavity and magnon modes, and $\Omega$ ($\omega_{d}$) is the strength (frequency) of parametric drive. In the rotating frame at $\omega_{d}/2$, the Hamiltonian in Eq. (14) becomes

### v2

@@ III Nonreciprocal superradiant quantum phase transition
Figure 4: Bidirectional contrast ratio $\mathcal{I}$ as a function of the normalized drive strength $\Omega/\kappa_{a}$ and the detuning ratio $\Delta_{m}/\Delta_{a}$. The white background in the upper-right corner corresponds to the unstable region in Fig. 2(b). The vertical white solid line marks the critical drive $\Omega=\Omega_{1}$, and the black dashed curve indicates $\Omega=\Omega_{2}$. Other parameters are the same as those in Fig. 2.
Furthermore, we introduce the bidirectional contrast ratio:
which satisfies $0\leq\mathcal{I}\leq 1$. This parameter quantitatively characterize the degree of nonreciprocity in the SQPT: a nonzero $\mathcal{I}$ signals the emergence of nonreciprocal behavior. The minimum value $\mathcal{I}_{\rm min}=0$ corresponds to a reciprocal SQPT, while the maximum value $\mathcal{I}_{\rm max}=1$ represents an ideal nonreciprocal SQPT. In Fig. 4, we show the bidirectional contrast ratio $\mathcal{I}$ as a function of the normalized drive strength $\Omega/\kappa_{a}$ and the detuning ratio $\Delta_{m}/\Delta_{a}$. The white region in the upper-right corner corresponds to the unstable regime identified in Fig. 2(b). The vertical white solid line (black dashed curve) marks the critical drive strength $\Omega=\Omega_{1}=2.025\kappa_{a}$ ($\Omega=\Omega_{2}$). The bidirectional contrast ratio $\mathcal{I}$ attains its minimum value $\mathcal{I}_{\rm min}=0$ in the blue region ($\Omega<\Omega_{1}$), where the system remains in the normal phase for both $K>0$ and $K<0$, indicating a reciprocal SQPT. Notably, two red regions appear for $\Omega_{1}<\Omega<\Omega_{2}$, where $\mathcal{I}$ reaches its maximum value $\mathcal{I}_{\rm max}=1$, corresponding to an ideal nonreciprocal SQPT. In the lower (upper) red region, the system is in the superradiant phase (normal phase) for $K>0$ but remains in the normal phase (superradiant phase) for $K<0$. In the remaining region ($\Omega>\Omega_{2}$), the bidirectional contrast ratio lies between 0 and 1 (i.e., $0<\mathcal{I}<1$). Here, the system is in the superradiant phase for both $K>0$ and $K<0$, but the order parameters differ: $|M|^{2}(K>0)=|M|^{2}_{+}\neq|M|^{2}(K<0)=|M|^{2}_{-}$.
IV Discussions and conclusions
For implementing the proposed scheme, we consider a cavity magnonic system comprising a parametrically driven coplanar waveguide resonator (CWR) coupled to a YIG sphere. The parametric drive on the CWR is supplied by a flux-driven Josephson parametric amplifier (JPA). Experimentally, strong coupling between the CWR mode and the magnon mode in a YIG sphere has been demonstrated in Refs. Morris17 ; Song25 . Specifically, a coupling strength of $g_{m}/2\pi=8.17$ MHz was achieved at temperatures on the order of 10 mK Morris17 . In these setups, the frequencies of CWR and magnon modes are typically around 5 GHz, while their decay rates are on the order of 1 MHz Morris17 ; Song25 . Moreover, the magnon frequency can be readily tuned by adjusting the bias magnetic field. As demonstrated in Ref. Wang18 , the bistability of cavity magnon polaritons induced by MKE has been observed in the two cases of aligning the bias magnetic field along the [100] and [110] crystallographic axes of the YIG sphere, corresponding to positive and negative MKE coefficients, respectively. The action of the flux-driven JPA on the CWR is well described by a parametric drive Hamiltonian [cf. the last term in Eq. (14)] Yamamoto08 ; Zhong13 , where the drive strength (e.g., tunable from 0 to 6 MHz Krantz16 ) and frequency can be precisely controlled via the external flux drive applied to the JPA. Given these experimentally accessible parameters and capabilities, the proposed scheme can be realistically realized in a CWR-magnon hybrid system.
In summary, we have demonstrated that the MKE in a cavity magnonic system offers a viable pathway to achieve a nonreciprocal SQPT Chiacchio23 ; Zhu24 ; Xu24 . The proposed setup, comprising a YIG sphere coupled to a microwave cavity, exhibits markedly different phase diagrams under positive and negative MKE coefficients—controlled by orienting the bias magnetic field along different crystallographic axes of the YIG sphere Wang18 ; Zhang19 . These distinct phase boundaries directly lead to nonreciprocal critical behavior, as confirmed by analyzing the magnon occupation and its fluctuations around the transition. The introduction of a bidirectional contrast ratio further allows quantitative characterization of the nonreciprocity. Our results establish cavity magnonics as a promising platform for exploring nonreciprocal SQPT and may inform future designs of quantum devices requiring directional control.
@@ Acknowledgments
This work is supported by the National Natural Science Foundation of China (Grants No. 12205069 and No. 12204139), the HZNU scientific research and innovation team project (Grant No. TD2025003), and the Hangzhou Leading Youth Innovation and Entrepreneurship Team project (Grant No. TD2024005). Y.H.K. is supported by the National Key Research and Development Program of China (Grant No. 2024YFA1408900). W.X. is supported by the Natural Science Foundation of Zhejiang Province (Grant No. LY24A040004), the Zhejiang Province Key R&D Program of China (Grant No. 2025C01028), and the Shenzhen International Quantum Academy (Grant No. SIQA2024KFKT010).
@@
where $a^{{\dagger}}$ and $m^{{\dagger}}$ ($a$ and $m$) are the creation (annihilation) operators for the cavity and magnon modes with frequencies $\omega_{a}$ and $\omega_{m}$, respectively, $K$ is the MKE coefficient, $g_{m}$ is the coupling strength between cavity and magnon modes, and $\Omega$ ($\omega_{d}$) is the strength (frequency) of parametric drive. In the rotating frame at $\omega_{d}/2$, the Hamiltonian in Eq. (14) becomes